OpenStudy (anonymous):

Evaluate the integral \[\int\limits_{}^{}( x/\sqrt{x^2-3x+10})dx\]

OpenStudy (mr.math):

Use a trigonometric substitution after completing the square of the expression under the radical.

OpenStudy (anonymous):

Write x as m(2x-3)+n m=1/2 and n=3/2. so spererating the terms you get \[(1/2) \int\limits_{}^{}(2x-3)/\sqrt{x ^{2}-3x+10} + \] the second term which is \[3/2\int\limits_{}^{}1/\sqrt{x ^{2}+3x+10}\]

OpenStudy (anonymous):

Now both are integrable, the first one is of the form f'(x)/f(x) hence after integration it becomes 2 sqrt(x^2+3x+10). The second term complete the square and write denominator as (x+3/2)^2+10-9/4 And then apply the standard formula.